OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian? A. 36 wraps B. 37 wraps C. 30 wraps D. 34 wraps

OpenStudy (anonymous):

Let the number of vegetarian wraps be v and the number of chicken wraps be c. v+c=70-(1) 1.0v+1.8c=98.8 10v+18c=988-(2) (1)*10=>10v+10c=700-(3) (2)-(3)=>8c=288 c=288/8 c=36 v+36=70 v=70-36 =34